OpenStudy (anonymous):

Jerry’s friend Brandon is planning to start saving for a trip at the same time as Jerry. The graph shows the balance in Brandon’s account over the first few months of saving. Whose account will have the greatest balance after the first five months of savings, Brandon’s or Jerry’s?

OpenStudy (ltrout):

Post the graph?

OpenStudy (thewafflebro):

any graph?

OpenStudy (thewafflebro):

what do you think

OpenStudy (anonymous):

What do I think? I have no clue

OpenStudy (thewafflebro):

dude you are dumb i cant just tell you the answer this is to help dude just have a best guess at leas goddam

OpenStudy (thebossofme191):

ooooo

OpenStudy (thebossofme191):

calm down dude

OpenStudy (anonymous):

Yeah seriously

OpenStudy (thewafflebro):

its your fault youre just coming for the answer dude at least have a guess

OpenStudy (anonymous):

Right but you don't have to be rude, have some respect for other people...

OpenStudy (thebossofme191):

ight both ya'll shut yo faces and i'll help u figure dis sht out

OpenStudy (anonymous):

Thanks Thebossofme191

OpenStudy (thebossofme191):

np

OpenStudy (thewafflebro):

why for cheating oh yeah ill respect you dude for cheating CHEATOR

OpenStudy (anonymous):

😂

OpenStudy (thebossofme191):

bro chill

OpenStudy (thewafflebro):

i will if he doesint chill

OpenStudy (thebossofme191):

rephrase that plz

OpenStudy (thebossofme191):

ight look at the graph for what is it 5 weeks or months

OpenStudy (anonymous):

Its 5 months

OpenStudy (thebossofme191):

ahh 5 months

OpenStudy (thewafflebro):

TROLOLOLOLOLOLO YOUVE BEEN TROLLED SAY HI TO YOUTUBE

OpenStudy (thewafflebro):

HAHAHAHAHAAHHAHAHAHAHAHHAHAHHAHAHHAHAHAHHAHAAAHA

OpenStudy (thebossofme191):

and tell me how much Brandon has

OpenStudy (thewafflebro):

YOU TWO REALLY THOUGHT I WAS A HATER LOL

OpenStudy (anonymous):

Brandon starts with $0

OpenStudy (anonymous):

Yeah as you can see neither of us really care so bye...

OpenStudy (thewafflebro):

LOL GOOD ONE

OpenStudy (anonymous):

Lol yeah bye

OpenStudy (thebossofme191):

yes he starts with $0 but where is he at after 5months

OpenStudy (anonymous):

I think $350 but I'm not positive

OpenStudy (thebossofme191):

i would say around there

OpenStudy (anonymous):

So now we need to know where Jerry will be after 5 months

OpenStudy (thebossofme191):

do u have a graph for Jerry

OpenStudy (anonymous):

No but I know that this is the formula. f(t) = 2^t-1

OpenStudy (thebossofme191):

ahh

OpenStudy (anonymous):

Does that help?

OpenStudy (thebossofme191):

can u tell me what f equals

OpenStudy (anonymous):

It doesn't tell me what f equals, but t = months

OpenStudy (thewafflebro):

BUT ACCTULLY DONT CHEAT BRO

OpenStudy (thebossofme191):

ok so f(months)=2^months-1

OpenStudy (anonymous):

Okay, so how do I find Jerry's total after 5 months

OpenStudy (thebossofme191):

2^5-1

OpenStudy (thebossofme191):

2^5-1=?

OpenStudy (anonymous):

2^5-1 is 2^4 so 2^4 = 16

OpenStudy (thebossofme191):

try 2^5=25-1=24

OpenStudy (anonymous):

Ahhh I see, order of operation is important

OpenStudy (thebossofme191):

PEMDAS

OpenStudy (thebossofme191):

Exponents first subtract second

OpenStudy (anonymous):

So, Jerry would have how much in his account after 5 months? $240? Just a guess

OpenStudy (thebossofme191):

i think so sry i didnt respond my wifi went out

OpenStudy (thebossofme191):

@04217_telstar

OpenStudy (anonymous):

Thats okay

OpenStudy (anonymous):

Thank you for your help

OpenStudy (thebossofme191):

np